OpenStudy (anonymous):

Evaluate the derivative using logarithmic differentiation if y = x(x + 5)^3/(6x − 1)^2

OpenStudy (bahrom7893):

Lny = Ln[x(x + 5)^3/(6x − 1)^2] Lny = Lnx + Ln(x+5)^3 - Ln(6x-1)^2

OpenStudy (bahrom7893):

Lny = Lnx + 3 Ln(x+5) - 2 Ln(6x-1) (1/y)y' = (1/x) + 3/(x+5) - 2/(6x-1)

OpenStudy (bahrom7893):

y' = y*[(1/x) + 3/(x+5) - 2/(6x-1)]

OpenStudy (bahrom7893):

and now just replace y with the equation taht u had for y

OpenStudy (anonymous):

the original one at the very top?

OpenStudy (paxpolaris):

Yes \[ y = {x(x + 5)^3 \over (6x − 1)^2}\]

OpenStudy (anonymous):

its saying its wrong idk y

OpenStudy (anonymous):

http://dl.dropbox.com/u/14986544/photo.jpg

OpenStudy (anonymous):

so its -12?